Evaluate det(−A) and det(−AT). Compare det(−A) and det(−AT) for:


(2.1) A = −4 2

3 −3 ;



(2.2) A = 3 1 −2

−5 3 −6

−1 0 −4


1
Expert's answer
2021-05-31T19:02:12-0400

2.1)


det( A ) = "\begin{vmatrix}
 -4 & 2 \\
 3 & -3 \\
\end{vmatrix}"



∴ det( -A ) = "\begin{vmatrix}
 4 & -2 \\
 -3 & 3
\end{vmatrix}"




Expanding along 1st row, we have


det(-A) = 4 * 3 - (- 3 ) * (- 2 )


det( -A ) = 6



Again, det( - AT) = "\begin{vmatrix}
 4 & -3 \\
 -2 & 3 \\
\end{vmatrix}"


Expanding along 1st row, we have


det(-A) = 4 * 3 - (- 2 ) * (- 3 )


det(-AT) = 6



Hence, we see that det( -A ) = det(-AT) .




2.2)



det(A) = "\begin{vmatrix}
 3 & 1 & -2 \\
 -5 & 3 & -6 \\
-1&0&-4\\
\end{vmatrix}"


So, det( - A ) = "\begin{vmatrix}
 -3 & -1 & 2 \\
 5 & -3 & 6 \\
1&0&4\\
\end{vmatrix}"


Expanding along 1st row, we have


det(-A) = -3[ ( -3 ) "*" 4 - 0 "*" 6] - ( -1 )[5 "*" 4 - 1 "*" 6] + 2[ 5 "*" 0 - 1 "*" (-3)]


det(-A) = 56




det(- AT) = "\begin{vmatrix}
 -3 & 5 & 1 \\
 -1 & -3 & 0 \\
2&6&4\\
\end{vmatrix}"


Expanding along 1st row, we have


det(-A) = -3[ ( -3 ) "*" 4 - 6 "*" 0] - 5 [ ( -1 ) "*" 4 - 2 "*" 0] + 1[ ( -1 ) "*" 6 - 2 "*" ( -3 ) ]



det(- AT) = 56


Hence, we see that det( -A ) = det(-AT) .